OpenStudy (anonymous):

Find the tangent line of the curve defined by the vector equation at the point t=0. u(t)=(1+t^3) i+te^(-t) j+e^(2t) k

OpenStudy (amistre64):

the tangent line of a vector defines function is the derivatives of each component

OpenStudy (amistre64):

<3t^2, -t e^-t + e^-t, 2e^2t> if i see it right

OpenStudy (amistre64):

at t=0 we got: <0,0,2> right? plug that into the point to get the line

OpenStudy (amistre64):

x = 1+t^3 ; x' = 3t^2 y = t e^(-t) ; y' = -t e^(-t) + e^(-t) z = e^(2t) ; z' = 2 e^(2t) at t=0 we get the point (1,0,1) the tangent vector at t=0 is: n<0,1,2>

OpenStudy (amistre64):

the line equation in then: x=1 y=1n z=1 +2n where n is the scalar for the tangent vector

OpenStudy (amistre64):

do youagree?

OpenStudy (anonymous):

but the point should be <0,0,1>

OpenStudy (amistre64):

hmm; x = 1+t^3 = 1 + (0) =1 y = t e^(-t) = 0(e^...) = 0 z = e^(2t) = e^0 = 1 point (1,0,1) ; how do we get x=o?

OpenStudy (anonymous):

ohhh, you are right